Recent Log Entries
iocsoftfilter restored
iocsoftserptn05 restored
Girder 3 vacuum work complete, ion pump is ON
We spent a little time getting girder3 coarse aligned, things look good now. Then Phil added the harp and blankoffs here and there, tightening all flanges. We pumped with turbo a few hours, the ion pump is now ON at ~ 300uA and the pump cart was valved OUT (RA valve to pump cart closed).
Once the MBO magnet arrives, S&A will do a 2A alignment of the dipole and girder3. Nice.
hipotted the NEG pump on the DP can phase2 beamline
I used a gamma vacuum supply to hipot the ion pump hanging from the dif pump can on the phase2 beamline: VIP0I06B. This seemed to blow off an emitter. Ion pump current now relatively stable at ~ 200nA. The other beamline ion pump in this region, VIP0I07, near the NEG pump Marcy activated yesterday, might also have a field emitter. This pump current is about 1 uA. Will investigate tomorrow.
NEG activation VGP0I06 and VGP1D00A
Yesterday we conditioned the two phase2 NEG pumps mentioned the log entry title, today we performed the full activation.
Conditioning (yesterday, p.m.): 3 V, 15 minute ramp, 30 minute soak
Activation (today, a.m.): 8 V, 15 minute ramp, 60 minute soak
For these tasks, the beamline ion pumps were OFF, with gas expelled from beamline into pump cart. Pressure at the pump cart increased as expected, and then 'rolled over' indicating successful activation. Some hours after the activation, the nearby beamline ion pumps are now ON and reading ~ 1 uA each.